# TinyML NLP Scheme for Semantic Wireless Sentiment Classification with Privacy Preservation

### v2

@@ Abstract
Natural Language Processing (NLP) operations, such as semantic sentiment analysis and text synthesis, often raise privacy concerns and demand significant on-device computational resources. Centralized learning (CL) on the edge provides an energy-efficient alternative but requires collecting raw data, compromising user privacy. While federated learning (FL) enhances privacy, it imposes high computational energy demands on resource-constrained devices. This study provides insights into deploying privacy-preserving, energy-efficient NLP models on edge devices. We introduce semantic split learning (SL) as an energy-efficient, privacy-preserving tiny machine learning (TinyML) framework and compare it to FL and CL in the presence of Rayleigh fading and additive noise. Our results show that SL significantly reduces computational power and CO2 emissions while enhancing privacy, as evidenced by a fourfold increase in reconstruction error compared to FL and nearly eighteen times that of CL. In contrast, FL offers a balanced trade-off between privacy and efficiency. This study provides insights into deploying privacy-preserving, energy-efficient NLP models on edge devices.
@@ I Introduction
Artificial Intelligence (AI) has gained widespread adoption, with applications ranging from text and image classification to text and image generation. NLP has seen rapid growth, driving advancements in virtual assistants and language classification systems. Transformer-based Large Language Models (LLMs), such as OpenAI’s GPT series [1], and BERT, have significantly advanced NLP by enabling machines to perform complex tasks with high accuracy. However, these models require substantial computational resources for training and inference, making them less practical for deployment on resource-constrained devices. Additionally, older temporal models like long short-term memory networks (LSTMs) and other recurrent neural network (RNN) variants remain a popular choice due to their ability to balance computational efficiency and performance.
Firstly, although more efficient than LLMs, temporal models still require substantial storage and processing power. The high dimensionality of language data and large vocabularies further amplify the computational demands [2]. Secondly, privacy remains a major concern, as training robust models often necessitates diverse datasets, potentially exposing sensitive information. CL exacerbates this risk by collecting raw data, whereas FL mitigates it by decentralizing training. However, FL remains vulnerable to inference attacks on gradient updates. SL enhances privacy by transmitting only intermediate activations, significantly reducing the risk of data reconstruction. Thirdly, deploying models on edge devices introduces challenges, including limited computational power, bandwidth constraints, and latency issues. Wireless data transmission over channels such as WiFi is particularly susceptible to noise, fading, and unstable connections, all of which can degrade communication efficiency.
@@ II-B Split Learning System
where $W_{\text{user}}^{(k)}$ are the user-side model weights at cycle $k$, and $S_{i}^{(k)}$ represents the output of the user’s model partition (i.e, smashed data). The output $S_{i}^{(k)}$ is first encoded into bits $X_{i}^{(k)}$ and then digitally modulated into a signal $Z_{i}^{(k)}$. This signal is transmitted through the wireless channel as will be defined in Section II-C. Due to channel effects, errors can be introduced in the received signal, affecting the accuracy of the demodulated bit stream $\hat{X}_{i}^{(k)}$. Upon receiving the noisy signal, the server demodulates and decodes it to obtain the estimated activations $\hat{S}_{i}^{(k)}$, which are then utilized to complete the forward pass
@@ II-D Communication Energy Calculation
In our approach, communication energy is determined by calculating the channel capacity $C$, which represents the maximum data rate for error-free transmission. Using the Shannon-Hartley theorem [14], the channel capacity incorporates the Signal-to-Noise Ratio (SNR), multiplied with the Rayleigh fading $f$ and the bandwidth $B$. The time required to transmit one bit and the corresponding energy consumption per bit is derived from the following equations. The SNR represents the ratio of transmission power $P$ to noise power $\sigma^{2}$. After calculating the SNR we proceed to calculate the channel capacity $C$ as
@@ III Experiments and Results
We conducted our experiments using the Sentiment140 dataset [15], which contains 1.6 million tweets. To adapt the dataset for resource-constrained devices, only the text and target labels (0 for negative sentiment and 1 for positive sentiment) were retained, and the dataset size was halved. All experimental configurations are summarized in Table I. Binary phase shift keying (BPSK) is the digital modulation scheme adopted to transmit the quantized data. Energy consumption and CO2 emissions were monitored using the Eco2AI framework [16], with measurements taken every 10 seconds during the FL cycles.
In the CL setup, 3 users collaborated in sending the data to the server, and the SL setup involved one user and one server. Both CL and FL models were trained for 50 cycles using the same batch size across all experiments. In SL, the model was partitioned at a specific split layer $L$, where convolutional and pooling layers were executed on the user’s device to minimize computational complexity, while the rest of the layers were handled by the server. The update rules for the SGD optimizer are defined as
@@ III-A2 Split learning
In our SL model, the user-side computation is represented by the function $f_{u}(x)$, where $x$ is the input data processed through the initial layers of the model, including convolutional and pooling layers. The output of $f_{u}(x)$ is then sent to encoder for compression then sent to the server, where the server-side computation $f_{s}(f_{u}(x))=y$ takes input and pass it to a decoder for decompression then processes the data further using the LSTM and subsequent layers to produce the final output $y$. This split allows the server to handle the more computationally intensive tasks.
@@ III-B Experimental Results
Our experiments are designed to evaluate and compare various learning frameworks based on three key aspects: model accuracy, computational energy consumption, and communication energy consumption. We also investigate the effects of noise and Rayleigh fading.
Fig. 3(a), shows that both CL and FL with 8-bit (Q8) and 32-bit (Q32) quantization converge at approximately 0.78 accuracy. However, introducing noise with an SNR level of 20 dB and fading in CL as shown in Fig. 3(d) causes a slight degradation in accuracy. This degradation occurs because CL transmits raw data, which is directly impacted by noise and fading during transmission. In contrast, FL transmits quantized model weights, which are smaller and well-structured, making them less vulnerable to transmission impairments. As shown in Fig. 3(b), lower quantization levels (e.g., Q4) has less accuracy due to precision loss, while Q8 and higher offer a good trade-off between accuracy and communication efficiency. Q8 stands out as the optimal choice, balancing both effectively.
@@
Table II elucidates a comprehensive comparison of the three approaches. The evaluation reveals several key insights about privacy-computation tradeoffs. The proposed SL scheme demonstrates the highest reconstruction error of 0.2681, indicating stronger privacy preservation as it becomes increasingly difficult to recover raw data from the intermediate activations $S_{i}$. In contrast, FL and CL show concerning privacy vulnerabilities with a low reconstruction error of 0.0671 and 0.0154, respectively.
Regarding user-side computational efficiency, SL and FL were evaluated at 20 dB SNR, while CL shows zero user-side computational energy in the table. Though not shown here, CL’s substantial server-side energy consumption should be considered for a complete efficiency assessment. SL requires the lowest computational power of 3.45 J compared to FL’s 60.82 J. It incurs the highest communication energy costs due to frequent transmission of intermediate activations, even with compression applied. FL, on the other hand, balances communication and computing efficiency, making it suitable for scenarios with limited communication resources. In both federated and split learning approaches, computational energy on the user side dominates over communication energy, resulting in generally lower energy consumption for the proposed TinyML SL on the user side. SL demonstrates superior user-side energy efficiency, producing 10 times lower CO2 emissions than FL and 20 times lower than CL, as measured using the Eco2AI library. These further strengthen the position of the proposed SL framework as an ideal choice for TinyML NLP classification problems.
@@ IV Conclusion
We explored TinyML approaches for semantic text sentiment classification via FL and SL. Our approaches are designed to be energy-efficient and privacy-preserving alternatives to CL. It is clear that quantization techniques such as 8-bit quantization appeared to be optimum in our scenario. SL excels in reducing user-side computation and CO2 emissions, while maintaining accuracy in noisy conditions but incurs higher communication energy. FL, especially with Q8 quantization, offers an optimal balance between computational efficiency, communication cost, and data privacy, making it ideal for decentralized environments with limited bandwidth. The proposed semantic TinyML SL and FL approaches are considered very efficient for tiny devices with limited energy resources, where SL is the most privacy-preserving, energy-efficient and environment-friendly. Future work could extend these TinyML schemes to LLMs and integrate differential privacy to further enhance communication efficiency and security.

### v3

@@ Abstract
Natural Language Processing (NLP) operations, such as semantic sentiment analysis and text synthesis, often raise privacy concerns and demand significant on-device computational resources. Centralized learning (CL) on the edge provides an energy-efficient alternative but requires collecting raw data, compromising user privacy. While federated learning (FL) enhances privacy, it imposes high computational energy demands on resource-constrained devices. This study provides insights into deploying privacy-preserving, energy-efficient NLP models on edge devices. We introduce semantic split learning (SL) as an energy-efficient, privacy-preserving tiny machine learning (TinyML) framework and compare it to FL and CL in the presence of Rayleigh fading and additive noise. Our results show that SL significantly reduces computational power and CO2 emissions while enhancing privacy, as evidenced by a fourfold increase in reconstruction error compared to FL and nearly eighteen times that of CL. In contrast, FL offers a balanced trade-off between privacy and efficiency. Our code is available for replication at our GitHub repository: https://github.com/AhmedRadwan02/TinyEco2AI-NLP.
@@ I Introduction
Artificial Intelligence (AI) has gained widespread adoption, with applications ranging from text and image classification to text and image generation. NLP has seen rapid growth, driving advancements in virtual assistants and language classification systems. Transformer-based Large Language Models (LLMs), such as OpenAI’s GPT series [1] and BERT, have significantly advanced NLP by enabling machines to perform complex tasks with high accuracy. However, these models require substantial computational resources for training and inference, making them less practical for deployment on resource-constrained devices. Additionally, older temporal models like long short-term memory networks (LSTMs) and other recurrent neural network (RNN) variants remain a popular choice due to their ability to balance computational efficiency and performance.
Although more efficient than LLMs, temporal models still require substantial storage and processing power. The high dimensionality of language data and large vocabularies further amplify the computational demands [2].
Privacy is another critical concern, as training robust models often necessitates diverse datasets, potentially exposing sensitive information. While CL exacerbates this risk by collecting raw data, FL mitigates it through decentralized training. However, FL remains vulnerable to inference attacks on gradient updates. SL addresses this vulnerability by transmitting only intermediate activations, significantly reducing the risk of data reconstruction.
In addition to privacy and computation, edge deployment introduces further challenges. Limited processing capacity, bandwidth constraints, and latency issues can hinder model performance. Moreover, wireless data transmission over channels such as WiFi is particularly susceptible to noise, fading, and unstable connections, which can degrade communication efficiency.
@@ II-B Split Learning System
where $W_{\text{user}}^{(k)}$ are the user-side model weights at cycle $k$, and $S_{i}^{(k)}$ represents the output of the user’s model partition (i.e., smashed data). The output $S_{i}^{(k)}$ is first encoded into bits $X_{i}^{(k)}$ and then digitally modulated into a signal $Z_{i}^{(k)}$. This signal is transmitted through the wireless channel as will be defined in Section II-C. Due to channel effects, errors can be introduced in the received signal, affecting the accuracy of the demodulated bit stream $\hat{X}_{i}^{(k)}$. Upon receiving the noisy signal, the server demodulates and decodes it to obtain the estimated activations $\hat{S}_{i}^{(k)}$, which are then utilized to complete the forward pass
@@ II-D Communication Energy Calculation
In our approach, communication energy is determined by calculating the channel capacity $C$, which represents the maximum data rate for error-free transmission. Using the Shannon-Hartley theorem [14], the channel capacity incorporates the Signal-to-Noise Ratio (SNR), multiplied by the Rayleigh fading $f$ and the bandwidth $B$. The time required to transmit one bit and the corresponding energy consumption per bit is derived from the following equations. The SNR represents the ratio of transmission power $P$ to noise power $\sigma^{2}$. After calculating the SNR we proceed to calculate the channel capacity $C$ as
@@ III Experiments and Results
We conducted our experiments using the Sentiment140 dataset [15], which contains 1.6 million tweets. To adapt the dataset for resource-constrained devices, only the text and target labels (0 for negative sentiment and 1 for positive sentiment) were retained, and the dataset size was halved. All experimental configurations are summarized in Table I. Binary Phase Shift Keying (BPSK) is the digital modulation scheme adopted to transmit the quantized data. Energy consumption and CO2 emissions were monitored using the Eco2AI framework [16], with measurements taken every 10 seconds during the FL cycles. In the CL setup, 3 users collaborated in sending the data to the server, and the SL setup involved one user and one server. Both CL and FL models were trained for 50 cycles using the same batch size across all experiments. In SL, the model was partitioned at a specific split layer $L$, where convolutional and pooling layers were executed on the user’s device to minimize computational complexity, while the rest of the layers were handled by the server. The update rules for the SGD optimizer are defined as
@@ III-A2 Split learning
In our SL model, the user-side computation is represented by the function $f_{u}(x)$, where $x$ is the input data processed through the initial layers of the model, including convolutional and pooling layers. The output of $f_{u}(x)$ is then sent to encoder for compression then sent to the server, where the server-side computation $f_{s}(f_{u}(x))=y$ takes input and passes it to a decoder for decompression then processes the data further using the LSTM and subsequent layers to produce the final output $y$. This split allows the server to handle the more computationally intensive tasks.
@@ III-B Experimental Results
We evaluate and compare learning frameworks based on model accuracy, computational energy, and communication energy while also examining the effects of noise and Rayleigh fading. Fig.3(a) shows that both CL and FL, with 8-bit (Q8) and 32-bit (Q32) quantization, converge at approximately 0.78 accuracy. However, introducing noise with an SNR level of 20 dB and Rayleigh fading in CL, as shown in Fig.3(d), leads to a slight degradation in accuracy. This degradation occurs because CL transmits raw data, which is directly affected by noise and fading during transmission. In contrast, FL transmits quantized model weights, which are smaller and more structured, making them less susceptible to transmission impairments. As shown in Fig. 3(b), lower quantization levels (e.g., Q4) reduce accuracy due to precision loss, while Q8 and higher offer a favorable trade-off between accuracy and communication efficiency. Q8 emerges as the optimal balance.
@@
Table II compares the three approaches. The evaluation reveals key insights about privacy-computation tradeoffs. The proposed SL scheme demonstrates the highest reconstruction error of 0.2681, indicating stronger privacy preservation as it becomes increasingly difficult to recover raw data from intermediate activations $S_{i}$. In contrast, FL and CL show concerning privacy vulnerabilities with low reconstruction errors of 0.0671 and 0.0154, respectively.
For user-side computational efficiency, SL and FL were evaluated at 20 dB SNR, while CL shows zero user-side energy in the table. However, CL’s substantial server-side energy—though not shown—should be considered. SL requires the least computational energy, reaching 3.45 J, compared to FL 60.82 J, but incurs the highest communication energy due to frequent intermediate activation transfers, even with compression. FL balances between communication and computation, making it suitable for communication-limited scenarios. In both FL and SL, user-side computation dominates energy use, with SL achieving the lowest overall user energy in TinyML settings. SL also offers superior efficiency, emitting 10 times less CO2 than FL and 20 times less than CL, as measured by Eco2AI. These results support SL as an ideal choice for TinyML NLP classification.
@@ IV Conclusion
We explored TinyML approaches for semantic text sentiment classification via FL and SL. Our approaches are designed to be energy-efficient and privacy-preserving alternatives to CL. It is clear that quantization techniques such as 8-bit quantization appeared to be optimum in our scenario. SL excels in reducing user-side computation and CO2 emissions while maintaining accuracy in noisy conditions but incurs higher communication energy. FL, especially with Q8 quantization, offers an optimal balance between computational efficiency, communication cost, and data privacy, making it ideal for decentralized environments with limited bandwidth. The proposed semantic TinyML SL and FL approaches are considered highly efficient for tiny devices with limited energy resources, where SL is the most privacy-preserving, energy-efficient and environment-friendly. Future work could extend these TinyML schemes to LLMs and integrate differential privacy to further enhance communication efficiency and security.